# SpiderGen:Towards Procedure Generation For Carbon Life Cycle Assessments with Generative AI

### v2

@@ 4 SpiderGen: Our Proposed Solution
Quantity: To gain a sufficient understanding of a product category, we must ask for the right number of products to use. In early experiments, we found that asking for too few products would result in PFG information that was over-fitted to a specific niche within a product category. However, querying for too many products may make the PFG over-generalized and therefore inaccurate.(2) Generate Sample Product Process Lists: To find common processes across a given product category, SpiderGen instructs an LLM to select diverse products that use a wide variety of materials and processes. For each product, SpiderGen queries an LLM for details about what components make up the product, and how these components are processed. We then query an LLM to find the processes involved in manufacturing and distributing the product. We prompt an LLM to list the process name, the life cycle phase it is involved in, and describe why this process is included. We use the LCA ISO standards to guide the LLM for process generation to cover various types of processes and ensure complete coverage of environmental impacts (EPD).
@@
(4) Generate the final PFG: Given the processes that form the vertices of $G_{pc}$, we now generate the edges of the graph to place these processes in the correct ordering. We establish two types of ordering: explicit ordering and implicit ordering. Implicit ordering is based on the life cycle phases associated with each process. For example, processes that are a part of the “upstream” phase will always appear before processes that are a part of the “core” phase. After creating these implicit orderings, we use an LLM to generate explicit orderings within processes with the same life cycle phase. For example, the LLM will reason that a process called “Mixing dough” will appear before “Forming noodles with dough”, where both are core processes. From our observation, modern LLMs can reason about explicit ordering quite well, and provide orderings that align with real process orderings. By using implicit ordering based on life cycle phases dictated by the taxonomy of LCA, SpiderGen is able to generate more accurate explicit orderings within a smaller set of processes that belong to the same life cycle phase. Using this method, we observe that SpiderGen is able to create process orders that are close to the correct PFGs.
@@ 5 Evaluation Methodology
LLMDirect: The LLMDirect method directly prompts the LLM for the PFG using the well-established method of Chain-of-Thought Reasoning (sahoo2025systematicsurveypromptengineering). With this method, we provide the LLM with step-by-step instructions on how to gather all the necessary processes to create the PFG, as well as how to create the graph using these steps. We also prompt it to provide a rationale for each step.
@@
where $n_{specific}$ is the number of ground-truth nodes that are in “specific” pairings, $n_{subprocess}$ is the number of ground-truth nodes that are in “subprocess” pairings, $n_{match}$ is the number of ground-truth nodes that are in “match” pairings, $n_{groundtruth}$ is the total number of ground truth nodes, $n_{wrong}$ is the number of generated nodes that are labeled as “wrong”, and $n_{generated}$ is the total number of generated nodes.We provide a set of ten manually evaluated [^3] We initially utilized LLM-as-a-Judge techniques to rapidly conduct multiple of these evaluations. However, we observed that even with the addition of example cases, and chain-of-thought prompting, LLM-as-a-Judge tended to estimate a higher recall than a human judge. Additional details of this analysis can be found in the full version of the paper (synergylabs_spidergen) analyses, and describe themes in the erroneous processes produced by SpiderGen in Table 1. We observe that the majority of errors produced by SpiderGen come from two sources: (1) misalignment of an appropriate “scope” for the processes that must be included (i.e, how many tangential or auxiliary processes should be included), and (2) extrapolation and assumptions about how a product might be used and what materials may be involved to create the product.
@@
Category ID Product Category PMI F1-Score
@@
Table 1: Results of our Qualitative Analysis for 10 products (Shown circled in Figure 3). We order the products based on complexity denoted by the number of nodes (in parenthesis). We observe higher F1 scores for products with higher PMI values in general. Note that PMI is the quantitative metric defined in Equation 4.
@@
Figure 4: Comparing the Normalized PMI Scores for 65 products with varying complexity as denoted by the number of nodes in the ground truth $G_{pc}$ graphs. SpiderGen has higher normalized PMI for simpler product categories (e.g. “Dairy Products”) and lower for more complicated ones (e.g. “Railways”). We also report the qualitative scores for 10 different product categories (circled nodes) in Table 1, and include labels for some of the product categories in Table 1 in this figure.
@@ 6 Evaluation Results
We present an evaluation of SpiderGen, highlighting key results. Additional evaluation results can be found in the extended version of our paper (synergylabs_spidergen).
@@ 7 Discussion and Conclusion
Limitations and Open Problems:

### v3

@@ 4 SpiderGen: Our Proposed Solution
Quantity: To gain a sufficient understanding of a product category, we must ask for the right number of products to use. In early experiments, we found that asking for too few products would result in PFG information that was over-fitted to a specific niche within a product category. However, querying for too many products may make the PFG over-generalized and therefore inaccurate.
(2) Generate Sample Product Process Lists: To find common processes across a given product category, SpiderGen instructs an LLM to select diverse products that use a wide variety of materials and processes. For each product, SpiderGen queries an LLM for details about what components make up the product, and how these components are processed. We then query an LLM to find the processes involved in manufacturing and distributing the product. We prompt an LLM to list the process name, the life cycle phase it is involved in, and describe why this process is included. We use the LCA ISO standards to guide the LLM for process generation to cover various types of processes and ensure complete coverage of environmental impacts (EPD).
@@
(4) Generate the final PFG: Given the processes that form the vertices of $G_{pc}$, we now generate the edges of the graph to place these processes in the correct ordering. We establish two types of ordering: explicit ordering, and implicit ordering. Implicit ordering is based on the life cycle phases associated with each process. For example, processes that are a part of the “upstream” phase will always appear before processes that are a part of the “core” phase. After creating these implicit orderings, we use an LLM to generate explicit orderings within processes with the same life cycle phase. For example, the LLM will reason that a process called “Mixing dough” will appear before “Forming noodles with dough”, where both are core processes. From our observation, modern LLMs can reason about explicit ordering quite well, and provide orderings that align with real process orderings. By using implicit ordering based on life cycle phases dictated by the taxonomy of LCA, SpiderGen is able to generate more accurate explicit orderings within a smaller set of processes that belong to the same life cycle phase. Using this method, we observe that SpiderGen is able to create process orders that are close to the correct PFGs.
@@ 5 Evaluation Methodology
LLMDirect: The LLMDirect method directly prompts the LLM for the PFG using the well-established method of Chain-of-Thought Reasoning (sahoo2025systematicsurveypromptengineering). With this method, we provide the LLM with step-by-step instructions on how to gather all the necessary processes to create the PFG, as well as how to create the graph using these steps. We also prompt it to provide a rationale for each step.We include our exact prompt for this method in the Technical Appendix.
@@
where $n_{specific}$ is the number of ground-truth nodes that are in “specific” pairings, $n_{subprocess}$ is the number of ground-truth nodes that are in “subprocess” pairings, $n_{match}$ is the number of ground-truth nodes that are in “match” pairings, $n_{groundtruth}$ is the total number of ground truth nodes, $n_{wrong}$ is the number of generated nodes that are labeled as “wrong”, and $n_{generated}$ is the total number of generated nodes.We provide a set of ten manually evaluated  [^3] We initially utilized LLM-as-a-Judge techniques to rapidly conduct multiple of these evaluations. However, we observed that even with the addition of example cases, and chain-of-thought prompting, LLM-as-a-Judge tended to estimate a higher recall than a human judge. Additional details of this analysis can be found in the technical appendix. analyses, and describe themes in the erroneous processes produced by SpiderGen in Table 1. We observe that the majority of errors produced by SpiderGen come from two sources: (1) misalignment of an appropriate “scope” for the processes that must be included (i.e, how many tangential or auxiliary processes should be included), and (2) extrapolation and assumptions about how a product might be used and what materials may be involved to create the product.
@@
ID Product Category PMI F1-Score
@@
Table 1: Results of our qualitative analysis for 10 products (Circled in Figure 4). We order the products based on the number of nodes (in parenthesis). Note that PMI is the quantitative metric defined in Equation 4.
@@
Figure 4: Comparing the Normalized PMI Scores for 65 products with varying complexity as denoted by the number of nodes in the ground truth $G_{pc}$ graphs. SpiderGen has higher normalized PMI for simpler product categories (e.g. “Dairy Products”) and lower for more complicated ones (e.g. “Railways”). We also report the qualitative scores for 10 product categories (circled nodes) in Table 1, and label some of the product categories in Table 1 in this figure.
@@ 6 Evaluation Results
We present an evaluation of SpiderGen, highlighting key results. Additional evaluation results can be found in the extended version of our paper (sitaraman2025spidergenproceduregenerationcarbon).
@@ 7 Discussion and Conclusion
Limitations and Open Problems
@@ Acknowledgments
Appendix A Technical Appendix
This section contains additional information about the SpiderGen workflow implementation, including prompts used for the workflow, as well as additional results.
Additional Details About the Product Category Rules Dataset
We provide additional details about the dataset of product category rules that we utilize for our evaluation in Table 4. We note that the product categories that we use in our dataset have a wide range of complexity, ranging from including just 12 processes to 46. We also note that we analyze products with varying levels of “specificity”, as indicated by the length of the shortest UN CPC code which corresponds to the product category, ranging from 2 digits (broadest product category) to 5 digits (most specific or niche product category).
Figure 5: An example comparison of PFGs between (a) SpiderGen, and two baselines, (b) LLMDirect and (c) LLMExample for the product category “Asphalt Mixtures”. We note that SpiderGen captures all downstream processes, and captures the majority of core processes, as compared to both LLMDirect and LLMExample, which fail to capture large portions of important processes, such as processing Asphalt additives, and maintaining asphalt.
Product Category PMI ROUGE-L BLEU BERTScore Precision Recall F1-Score LLM-as-a-Judge
Railways (44) 0.02 0.08 0.0 0.74 1 0.25 0.4 0.68
Shower Enclosures (40) 0.02 0.14 0.01 0.74 0.9 0.63 0.74 0.73
T-Shirts, Tops (28) 0.03 0.18 0.05 0.8 1 0.63 0.77 0.80
Moka Coffee (22) 0.02 0.18 0.07 0.72 0.7 0.64 0.66 0.65
Dairy Products (22) 0.05 0.17 0.06 0.84 0.59 0.68 0.63 0.65
Graphite Products (20) 0.04 0.16 0.1 0.82 0.63 0.61 0.62 0.84
Detergents & Washing (19) 0.09 0.15 0.05 0.73 0.6 0.57 0.59 0.55
Woven Fabric (17) 0.05 0.17 0.1 0.81 1 0.73 0.84 0.71
Air Ducts (17) 0.07 0.13 0.28 0.72 0.71 0.59 0.65 0.65
Bottled Water (12) 0.03 0.23 0.2 0.76 1 0.42 0.59 0.83
Table 3: Results of SpiderGen for for 10 products (Shown circled in Figure 3) across multiple different metrics. We order the products based on complexity denoted by the number of nodes (in parenthesis). The F1-score, precision and recall are all provided by the qualitative analysis measure provided in Section 4. We note that PMI is the most closely correlated to recall out of all evaluation metrics, which is why we utilize this metric throughout the paper. We additionally note that the LLM-as-a-Judge metric typically provides a higher score than the human-generated F1-score due to observing more of the generated processes matched to ground truth processes.
Additional Evaluation Results for SpiderGen on Ground Truth Data
In this section, we provide additional evaluations of SpiderGen on our ground-truth dataset of PFGs.
Comparing the PFGs between both baselines and SpiderGen
Figure 5 shows a comparison between SpiderGen, and two baselines, LLMDirect and LLMExample, which are described in Section 5. We note that in this case, LLMDirect and LLMExample miss significantly more processes compared to SpiderGen, which misses just four. We additionally note that SpiderGen primarily misses auxiliary processes that are not a part of the main procedures needed to create, use and dispose of asphalt.
Efficacy of SpiderGen in different life cycle phases:
We explore the efficacy of SpiderGen in generating processes from different life cycle stages (either upstream, core, or downstream). Figure 6 shows the PMI scores of 65 different SpiderGen-generated PFGs. We observe that SpiderGen demonstrates much lower interquartile ranges, as well as a lower median PMI score in upstream and core process determination than in downstream process determinations.
We believe that the higher median PMI score for downstream processes is due to the fact that, unlike upstream and core processes which may be more niche to a specific product category, many of these downstream processes are standard across many product categories. For example, the process ”disposal of product” or ”waste management of product” will appear in a majority of product categories. This means that it is much easier for SpiderGen to predict downstream processes than other types of processes. We observe that SpiderGen typically generates as many of these standard downstream processes as they are relevant to the product category. For example,SpiderGen typically assumes all possible cases for the disposal of a product (ex. recycling, disposal, consumption, etc.).
However, it is often the case that downstream processes are under-specified in real ground truth LCA documents. We observe that for a large portion of ground truth documents, the method of collecting downstream processes is either unclear or optional. This indicates that for specific product categories, there may be a higher level of uncertainty in downstream processes, and an LCA expert would be required to use their own interpretation of the document to conduct the LCA in this scenario. This increases the standard deviation of the PMI for downstream processes, as SpiderGen scores highly when compared to well-specified downstream process lists, and poorly when compared to optional or under specified downstream process lists.
Efficacy of SpiderGen on varying product specificity:
LCAs are conducted for product categories at varying levels of specificity. For example, there is a PCR for “Grain Mill Products”, which is a broad category, and “Disposable Surgical Drapes, Gowns, Air Suits and Face Masks”, which is a more niche category. Our aim in designing SpiderGen was to evaluate how well the system performs on product categories of differing specificity. In this evaluation, we define “specificity” using UN CPC codes, which are hierarchical codes. UN CPC codes have a larger number of digits (a maximum of 5) when the product category descriptor is more specific or niche, and a lower number of digits (a minimum of 2), when a product category descriptor is more broad. In our case, we measure the specificity of a ground-truth PCR to be the minimum number of digits of it’s UN CPC code list. Figure 7 shows a box plot graph which shows the normalized PMI for more and less specific products. Based on these values, we can observe that SpiderGen accommodates broad categories best, with higher median PMIs for less specific product categories. We note that in our dataset, there are significantly fewer product categories that are at either extreme (with a CPC code length of 2 or with a CPC code length of 5), with just 3 product categories having a CPC code length of 2 and just 10 having a CPC code length of 5.
Additional evaluation methods
We find that a core challenge of this work is to address how to adequately evaluate SpiderGen using evaluation metrics. In this section, we provide an evaluation of SpiderGen under additional evaluation metrics, and provide additional rationale for the utilization of PMI in the main body of the work over alternative metrics.
Comparing PMI with traditional metrics
We first compare our PMI metric with metrics traditionally used to capture similarity, such as ROUGE-L, BLEU, and BERTScore. These comparisons are shown in Table 3, We compare these scores to the qualitative evaluation scores (F1-Score, Precision and Recall), and find that PMI has the strongest correlation with recall out of all of the options for our dataset of 10 products.
Utilization of LLM-as-a-Judge
In this section, we utilize an LLM-as-a-Judge technique to provide an evaluation of the correctness of the processes generated by SpiderGen. We implement LLM-as-a-Judge by listing out all SpiderGen-generated processes and all ground-truth processes in an LLM prompt. We first ask the LLM to determine whether each process that is generated by SpiderGen is included in the ground truth. From this, we are able to calculate precision, where
$precision=\text{num\_included}/\text{total\_num\_generated\_processes}$ (5)
We then ask the LLM the reverse question of which processes in the ground truth are given by the SpiderGen processes. From this, we are able to calculate recall, where
$recall=\text{num\_included}/\text{total\_num\_ground\_truth\_processes}$ (6)
We term the number of calculate the LLM-as-a-Judge score as an F1-score between this recall and precision value, which is calculated the same as in Equation 1 in the main text.
The implementation details of our LLM-as-a-Judge are given in our GitHub repository (https://github.com/synergylabs/SpiderGen). We evaluate LLM-as-a-Judge across 10 products. We find that LLM-as-a-Judge routinely overestimates the recall in comparison to the human evaluator (in our dataset of 10 products, the LLM-as-a-Judge overestimated the F1-Score by 22.4%). We find that the LLM-as-a-Judge method also underestimates precision by 11.9% in this dataset of 10 products.
Figure 6: Normalized Mutual Information of SpiderGen results across different Life Cycle Phases. We observe the variance in performance is higher for ”Downstream” processes than ”Upstream” or ”Core”.
Figure 7: Normalized Mutual Information of SpiderGen results across different Product Specificity levels. We observe the standard deviation in performance is higher for broader product categories (such as those with a UN CPC code length of 4 or less), and is smaller for product categories that are the most specific (UN CPC code of 5).
Implementation Details
We implement SpiderGen in Python using, NumPy, Pandas, and the all-mpnet-base-v2 model from the HuggingFace Sentence Transformer Library  [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2#all-mpnet-base-v2, as well as the KMeans function from the scikit-learn library  [^6] Pedregosa, F.; Varoquaux, G.; Gramfort, A.; Michel, V.; Thirion, B.; Grisel, O.; Blondel, M.; Prettenhofer, P.; Weiss, R.; Dubourg, V.; Vanderplas, J.; Passos, A.; Cournapeau, D.; Brucher, M.; Perrot, M.; and Duchesnay, E. 2011. Scikit- learn: Machine Learning in Python. Journal of Machine Learning Research, 12: 2825–2830.. We utilize OpenAI’s API as well for making our api calls.
For evaluating SpiderGen, we utilize the Llama-8b-Instruct model  [^7] https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md through HuggingFace Hub. We utilize the weights of this model. For further details on the use of the LLama-8b-Instruct model to calculate the PMI of two texts, refer to (xu_benchmarking_2025).
All code except for OpenAI model calls were run locally on an Apple MacBook Pro, with an M3 Pro chip (11-core CPU, 14-core GPU), with 18 GB RAM. Further implementation details are provided in our GitHub repository (https://github.com/synergylabs/SpiderGen).
Product Category Number of Processes UN-CPC Code Length
GRAPHITE PRODUCTS 20 4
OTHER SPECIAL- AND GENERAL-PURPOSE MACHINERY… 21 3
DISPENSING SYSTEMS 20 4
DAIRY PRODUCTS 22 3
BEARINGS, BEARING UNITS AND PARTS THEREOF 21 4
ELECTRICAL MOTORS AND GENERATORS AND PARTS THEREOF… 18 5
SKI FOOTWEAR 15 4
MEAT OF MAMMALS (FRESH, CHILLED OR FROZEN) 24 4
MACHINE-TOOLS FOR DRILLING, BORING OR MILLING… 21 3
MULTI-PURPOSE FILMS 21 3
SEATS 16 4
WOVEN, KNITTED OR CROCHETED FABRICS (OF SYNTHETIC FIBRES) 17 3
PARRAFIN WAX 18 3
GRAIN MILL PRODUCTS 19 3
RUBBER ARTICLES FOR FOOTWEAR 19 4
MOKA COFFEE 22 5
FISH, OTHERWISE PREPARED OR PRESERVED… 21 4
HIGHWAYS, STREETS AND ROADS 39 4
INDUSTRIAL FURNACES AND OVENS 25 4
WOVEN FABRICS OF SILK AND SILK-LIKE FIBRES 33 4
AIR DUCTS, SUBSTANTIAL MATERIALS (NON-CONSTRUCTION PRODUCT) 17 4
VEGETABLE JUICE, PLANT MILK, PLANT MILK BASED PRODUCTS… 21 4
SAUCES, MIXED CONDIMENTS AND MUSTARD 19 5
CLOSABLE FLEXIBLE PLASTIC PACKAGING 18 4
HOME AND SOHO GATEWAYS 22 5
PLASTICS IN PRIMARY FORM 34 3
CORROSION PROTECTION OF FABRICATED STEEL PRODUCTS 27 5
SHOWER ENCLOSURES 40 5
ASPHALT MIXTURES 36 4
TROUSERS, SHORTS AND SLACKS AND SIMILAR GARMENTS 28 3
TEXTILE YARN AND THREAD … 42 3
FOOTWEAR 23 3
MEAT OF POULTRY (FRESH, FROZEN OR CHILLED) 21 4
RAILWAYS 44 5
PREFABRICATED BUILDINGS 46 3
T-SHIRTS, TOPS, SINGLETS AND OTHER VESTS 28 3
BASIC CHEMICALS 18 3
DETERGENTS AND WASHING PREPARATIONS 19 5
TEA 22 5
PARTS AND ACCESSORIES FOR COMPUTING MACHINES… 17 4
PASTA COOKED, STUFFED OR OTHERWISE PREPARED; COUSCOUS 23 4
PASSENGER CARS 15 5
FABRICATED STEEL PRODUCTS, EXCEPT CONSTRUCTION PRODUCTS… 13 3
DISPOSABLE SURGICAL DRAPES, GOWNS AND AIR SUITS 19 4
PREPARATIONS USED IN ANIMAL FEEDING FOR FOOD-PRODUCING ANIMALS 16 3
PORT OPERATION SERVICES 17 3
EVENTS AND TOURISM SERVICES 32 2
NONWOVEN WIPES 13 5
PULPS, WOODS OR OTHER FIBROUS CELLULOSIC MATERIALS 24 4
ARABLE AND VEGETABLE CROPS 17 2
TEXTILE MANUFACTURING SERVICES, NON-APPAREL FABRICS… 28 3
PRESERVES AND PREPARATIONS OF MEAT… 16 4
BAKERY PRODUCTS 21 3
APPAREL, EXCEPT FUR AND LEATHER APPAREL 27 3
BASIC ALUMINIUM PRODUCTS AND SPECIAL ALLOYS 18 4
SERVICE OF PROVIDING WASHED AND STERILIZED REUSABLE SURGICAL DRAPES… 29 4
BOTTLED WATERS, NOT SWEETENED OR FLAVOURED 12 4
RARE-EARTH CONCENTRATES, OXIDES, METALS, AND MAGNETS… 15 4
TRANSPORT SERVICES 26 4
SWEATERS, JERSEYS, PULLOVERS, CARDIGANS, FLEECES… 31 3
RAW SUGAR, REFINED SUGAR, AND MOLASSES 22 4
ESCALATORS AND MOVING WALKS 34 4
FURNITURE, EXCEPT SEATS AND MATTRESSES 17 4
PROCESSED LIQUID MILK AND CREAM 14 3
BIRDS’ EGGS IN SHELL, FRESH 22 2
Table 4: The full list of all product categories that were evaluated in this work (the product categories with in-depth evaluations shown are highlighted). The product categories have a wide range of the number of processes (ranging from 12 processes to 46), and with a wide spread of product specificity (as indicated by the length of the shortest UN-CPC code)
Prompts Used for Baselines
In this section, we provide the prompts used to produce our two baselines (LLMDirect and LLMExample). The prompt for LLMDirect takes the input of a product category name (“product_name”) and its corresponding description (“product_description”). This prompt is shown in Figure 8.
    """You are an expert in Life Cycle Assessment (LCA) and Product Category Rules (PCR).
    Your goal is to give the Upstream, Core, and Downstream processes for a specific product,
    define the edges that connect them in a directed process flow graph.
    The given product name is {product_name}
    The given product description is {product_description}
    Based on the product name and description, follow the instructions below to
    create the process flow graph.
    1. List any upstream processes that are involved with creating the product.
    Describe each process
    2. List any core processes that are involved with creating the product.
    Describe each process
    3. List any downstream processes that are involved with creating
    the product. Describe each process
    Based on the process descriptions above,
    carefully follow the instructions below to create the directed process flow graph:
    1. Order each of the processes
    by the sequence in which they are done and create edges between each of these processes.
    2. If one of the subprocesses described above is a subprocess of another process,
    indicate this by creating an edge from
    the subprocess to the process (ex. subprocess --> process).
    Indicate that it is a subprocess.
    Be sure to include steps between the upstream,
    core and downstream processeses (ex. a transformation
    activity between an upstream and core process)
    so that the graph is connected throughout.
    4. Give a single JSON based
    on the information given above in the exact format given below:
    "processes": {{
                    <process_name> : {{
                        "description": <process_description>,
                        "process_category": <list either upstream, core or downstream>,
                        "is_subprocess": <list either subprocess or process>,
                        "input_nodes": [ input_node_1, input_node_2 ...
                        "output_nodes": [output_node_1, output_node_2 ...],
                        reasoning:  < provide a detailed description of the rationale>
                     <process_name_2> : {{
                        "description": <component_description>,
                        "process_category": <list either upstream, core or downstream>,
                        "is_subprocess": <list either subprocess or process>,
                        "input_nodes": [ input_node_1, input_node_2 ...
                        "output_nodes": [output_node_1, output_node_2 ...],
                        reasoning:  < provide a detailed description of the rationale>
    Important Instructions:
    1. Ensure that the JSON format given is followed exactly.
    Do not follow any other JSON format
    2. Ensure your response includes a clear and concise breakdown of each process,
    using the information provided in the input JSON.
    3. Be sure that each  process is as detailed as possible
    5. Provide details of all assumptions made and rationale behind each determination."""
Figure 8: Prompt used for LLMDirect
The prompt for LLMExample takes the input of a product category name (“product_name”) and its corresponding description (“product_description”). It takes an additional input of an example PFG (in the case of our implementation, the additional example was the PFG for “Baked Goods”). This prompt is shown in Figure 9.
    """You are an expert in Life Cycle Assessment (LCA) and Product Category Rules (PCR).
    Your goal is to give the Upstream, Core, and Downstream processes for a specific product,
    define the edges that connect them in a directed process flow graph.
    The given product name is {product_name}
    The given product description is {product_description}
    Below is an example PCR. Generate your response such that it
    emulates the life cycle phases, and scopes of this document
    The document is the following, which describes the PCR for baked goods:
    {example}
    Based on the product name and description, follow the
    instructions below to create the process flow graph.
    1. List any upstream processes that are involved with
    creating the product.
    Describe each process
    2. List any core processes that are involved with creating the product.
    Describe each process
    3. List any downstream processes that are involved with creating the product.
    Describe each process
    Based on the process descriptions above, carefully follow the
    instructions below to create the directed process flow graph:
    1. Order each of the processes by the sequence in which they are done and
    create edges between each of
    these processes.
    2. If one of the subprocesses described above is a subprocess of another process,
    indicate this by creating an edge from the
    subprocess to the process (ex. subprocess --> process). Indicate that it is a subprocess.
    Be sure to include steps between the upstream, core and downstream processeses
    (ex. a transformation activity between an upstream and core process)
    so that the graph is connected throughout.
    4. Give a single JSON based on the information given above in the exact format given below:
    "processes": {{
                    <process_name> : {{
                        "description": <process_description>,
                        "process_category": <list either upstream, core or downstream>,
                        "is_subprocess": <list either subprocess or process>,
                        "input_nodes": [ input_node_1, input_node_2 ...
                        "output_nodes": [output_node_1, output_node_2 ...],
                        ‘‘reasoning":  < provide a detailed description of the rationale>
                     <process_name_2> : {{
                        "description": <component_description>,
                        "process_category": <list either upstream, core or downstream>,
                        "is_subprocess": <list either subprocess or process>,
                        "input_nodes": [ input_node_1, input_node_2 ...
                        "output_nodes": [output_node_1, output_node_2 ...],
                        ‘‘reasoning":  < provide a detailed description of the rationale>
    Important Instructions:
    1. Ensure that the JSON format given is followed exactly. Do not follow any other JSON format
    2. Ensure your response includes a clear and concise breakdown of each process,
    using the information provided in the input JSON.
    3. Be sure that each process is as detailed as possible
    5. Provide details of all assumptions made and rationale behind each determination."""
Figure 9: Prompt used for LLMExample